# Multifidelity linear regression for scientific machine learning from scarce data

### v1

@@ 3.3 Optimality analysis for control variate coefficients
This section proves the optimality of the control variate coefficient definitions in (17) and (18). To simplify the exposition, results are proved for the bi-fidelity case where $K=2$, but the generalization of the results to the general $K>2$ case is straightforward. Section 3.3.1 shows that (17) defines the optimal scalar $\alpha\in\mathbb{R}$ for the estimator (12), while Section 3.3.2 shows that (18) defines the optimal matrix $A\in\mathbb{R}^{d\times d}$ for the estimator (14).
3.3.1 Scalar control variate coefficient
This section considers the multifidelity linear regression coefficient estimator (12), where the expectation $C_{XY}$ is estimated by the estimator $\hat{C}_{XY}^{\alpha}$ defined in (11) which uses a scalar control variate coefficient $\alpha$.
We are interested in making our multifidelity learned model robust to random realizations of the training data: we thus begin by bounding the variance of learned model predictions, conditioned on the input $z$, in terms of the generalized variance of the estimator $\hat{C}_{XY}^{\alpha}$.
@@ 3.3.1 Scalar control variate coefficient
For $i=1,\ldots,d$, let $(\lambda_{i},v_{i})$ denote the eigenpairs of $\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{\alpha},\hat{C}_{XY}^{\alpha}]$. Then,
$\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}_{\rm MF}^{\alpha})|Z=z]=\sum_{i=1}^{d}\lambda_{i}c_{i}\leq\|c\|_{2}\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{\alpha},\hat{C}_{XY}^{\alpha}]),$ (19)
where $c=[c_{1},\ldots,c_{d}]^{\top}\in\mathbb{R}^{d}$ is a constant vector dependent on $z$ but independent of the estimator $\hat{\beta}_{\rm MF}^{\alpha}$.
@@
$\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}_{\rm MF}^{\alpha})|Z=z]$ $=\tilde{x}^{\top}\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{\alpha},\hat{C}_{XY}^{\alpha}]\tilde{x}=\sum_{i=1}^{d}\lambda_{i}\left<\tilde{x},v_{i}\right>^{2}.$
The equality in (19) follows from the assignment $c_{i}=\left<\tilde{x},v_{i}\right>^{2}$. Then, note that $\sum_{i=1}^{d}\lambda_{i}c_{i}\leq\|\lambda\|_{2}\|c\|_{2}\leq\|\lambda\|_{1}\|c\|_{2}$. The conclusion follows from the fact that $\|\lambda\|_{1}=\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{\alpha},\hat{C}_{XY}^{\alpha}])$. ∎
Because Lemma 3.2 provides an upper bound on the conditional variance of the learned model predictions that is linear in the trace of the covariance of the estimator $\hat{C}_{XY}^{\alpha}$, we seek an $\alpha$ that minimizes this quantity. Since $\hat{C}_{XY}^{\alpha}$ is a vector-valued multifidelity estimator in $\mathbb{R}^{d}$, we now prove the following general result for choosing a scalar control variate coefficient to minimize the generalized variance of a vector control variate estimator.
@@
Let $g^{(1)}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ and $g^{(2)}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ be vector-valued functions, and define the following multifidelity estimator for the vector mean of $g^{(1)}$:
$\hat{g}_{\rm MF}^{\alpha}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(1)}(z_{i})+\alpha\left(\frac{1}{m_{2}}\sum_{i=1}^{m_{2}}g^{(2)}(z_{i})-\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(2)}(z_{i})\right),$ (20)
where $\{z_{i}\}_{i=1}^{m_{2}}$ are i.i.d. inputs as before. Let $C_{ij}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}=\operatorname{\mathbb{C}ov}[g^{(i)}(Z),g^{(j)}(Z)]$. Then, the autocovariance of the vector-valued multifidelity estimator is given by
$\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{\alpha},\hat{g}_{\rm MF}^{\alpha}]$ $=\frac{1}{m_{1}}C_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(\alpha^{2}C_{22}-\alpha C_{12}-\alpha C_{21}\right),$ (21)
and $\alpha^{*}=\arg\min_{\alpha\in\mathbb{R}}\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{\alpha},\hat{g}_{\rm MF}^{\alpha}])=\mathsf{Tr}(C_{12})/\mathsf{Tr}(C_{22})$.
@@
Expression (21) can be obtained by taking the covariance of (20), using the bilinearity of the covariance operator and the independence of the inputs $z_{1},\ldots,z_{m_{2}}$. The trace of the covariance of the multifidelity estimator is given by (recall $\mathsf{Tr}(M)=\mathsf{Tr}(M^{\top})$):
$\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{\alpha},\hat{g}_{\rm MF}^{\alpha}])=\frac{1}{m_{1}}\mathsf{Tr}(C_{11})+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(\alpha^{2}\mathsf{Tr}(C_{22})-2\alpha\mathsf{Tr}(C_{12})\right).$
@@
$2\alpha\mathsf{Tr}(C_{22})=2\mathsf{Tr}(C_{21})\implies\alpha=\mathsf{Tr}(C_{21})/\mathsf{Tr}(C_{22}).$
We note that the definition (12) of $\hat{C}_{XY}^{\alpha}$ has the form of (20) with $g^{(j)}(z)=x(z)f^{(j)}(z)$. We therefore use Lemma 3.3 with this choice of $g^{(j)}(z)$ to determine the $\alpha$ that minimizes the generalized variance of $\hat{C}_{XY}^{\alpha}$, which appears in Lemma 3.2.
@@
Let $\alpha$ be given by Lemma 3.3 applied to the functions $g^{(1)}(z)=x(z)f^{(1)}(z)$ and let $g^{(2)}(z)=x(z)f^{(2)}(z)$. Then,
@@
    the trace of the covariance of the estimator $\hat{C}_{XY}^{\alpha}$ is minimized, and
@@
3.3.2 Matrix coefficient analysis
This section considers the multifidelity linear regression coefficient estimator (14), where the expectation $C_{XY}$ is estimated by $\hat{C}_{XY}^{A}$ defined in (13). We follow a similar line of reasoning to Section 3.3.1, and first express the variance of learned model predictions, conditioned on the input $z$, in terms of the generalized variance of the estimate $\hat{C}_{XY}^{A}$.
Lemma 3.5.
For $i=1,\ldots,d$, let $(\lambda_{i},v_{i})$ denote the eigenpairs of $Cov[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}]$. Then,
$\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}_{\rm MF}^{A})|Z=z]=\sum_{i=1}^{d}\lambda_{i}c_{i}$ (22)
where $c_{1},\ldots,c_{d}$ are constants dependent on $z$ but independent of the estimator $\hat{\beta}_{\rm MF}^{A}$.
This follows directly from the proof of Lemma 3.2. In fact, we can show an upper bound on (22) that is analogous to the upper bound in (19). However, as we will now prove, with a matrix control variate coefficient $A\in\mathbb{R}^{d\times d}$ we can minimize the eigenvalues in (22) directly (rather than merely minimizing their sum). We show this result for a general vector multifidelity estimator.
Lemma 3.6.
Let $g^{(1)}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ and $g^{(2)}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ be vector-valued functions, and define the following multifidelity estimator for the vector mean of $g^{(1)}$:
$\hat{g}_{\rm MF}^{A}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(1)}(z_{i})+A\left(\frac{1}{m_{2}}\sum_{i=1}^{m_{2}}g^{(2)}(z_{i})-\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(2)}(z_{i})\right),$ (23)
where $A\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ and $\{z_{i}\}_{i=1}^{m_{2}}$ are i.i.d. inputs as before. Let $C_{ij}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}=\operatorname{\mathbb{C}ov}[g^{(i)}(Z),g^{(j)}(Z)]$. Then, the autocovariance of the vector-valued multifidelity estimator is given by
$\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A},\hat{g}_{\rm MF}^{A}]$ $=\frac{1}{m_{1}}C_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(AC_{22}A^{\top}-C_{12}A^{\top}-AC_{21}\right),$ (24)
and $A^{*}=\arg\min_{A\in\mathbb{R}^{d\times d}}\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A},\hat{g}_{\rm MF}^{A}])=C_{12}C_{22}^{-1}$.
Proof.
Let $D=A-C_{12}C_{22}^{-1}$. Then, we can rewrite (24) as
$\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A},\hat{g}_{\rm MF}^{A}]$ $=\frac{1}{m_{1}}C_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(D(C_{22})^{-1}D^{\top}-C_{12}(C_{22})^{-1}C_{21}\right)$
$=\left(\frac{1}{m_{1}}C_{11}-\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)C_{12}(C_{22})^{-1}C_{21}\right)+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)D(C_{22})^{-1}D^{\top}$
$\equiv S_{1}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)D(C_{22})^{-1}D^{\top}.$
Both terms in the above expression are symmetric, and both are positive semidefinite: the second term because it has the form $BB^{\top}$ and the first term because grouping $1/{m_{1}}$ terms yields the conditional covariance of $g^{(1)}$ given $g^{(2)}$ which must be positive semi-definite and the $1/{m_{2}}$ term is positive. Invoking Weyl’s inequality on the eigenvalues of Hermitian matrices (see [27, Corollary 4.3.12]), we have
$\lambda_{i}(S_{1})\leq\lambda_{i}\left(S_{1}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)D(C_{22})^{-1}D^{\top}\right),\quad\text{for }i=1,2,\ldots,d,$
where $\lambda_{i}(X)$ denotes the $i$-th eigenvalue of the Hermitian matrix $X$. Thus, each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A},\hat{g}_{\rm MF}^{A}]$ is minimized for the choice $D=0$, and the conclusion follows. ∎
The following corollary is a consequence of the proof of Lemma 3.6.
Corollary 3.7.
The choice of $A^{*}$ in Lemma 3.6 minimizes each eigenvalue of the covariance of the multifidelity estimator $\hat{g}_{\rm MF}^{A}$. That is,
$\lambda_{i}(\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A^{*}},\hat{g}_{\rm MF}^{A^{*}}])\leq\lambda_{i}(\operatorname{\mathbb{C}ov}[\hat{g}_{\rm MF}^{A},\hat{g}_{\rm MF}^{A}]),\qquad\text{for all }A\in\mathbb{R}^{d\times d}.$
Corollary 3.7 is a stronger result than Lemma 3.6, and enables us to show that this choice of $A^{*}$ minimizes the conditional variance of the multifidelity learned model predictions:
Theorem 3.8.
Let $A^{*}$ be given by Lemma 3.6 applied to the functions $g^{(1)}(z)=zf^{(1)}(z)$ and $g^{(2)}=zf^{(2)}(z)$. Then,
  1. 1.
    the generalized variance of the estimator $\hat{C}_{XY}^{A}$ is minimized,
  2. 2.
    the generalized variance of the estimated regression coefficients $\hat{\beta}_{\rm MF}^{A}$ is also minimized, and
  3. 3.
Proof.
The first statement follows directly from Lemma 3.6. The third statement follows from the Corollary 3.7 and Lemma 3.5. To show the second statement, note that
$\operatorname{\mathbb{C}ov}[\hat{\beta}_{\rm MF}^{A},\hat{\beta}_{\rm MF}^{A}]=(C_{XX})^{-1}\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}](C_{XX})^{-\top},$
so if $\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}]$ has eigendecomposition $\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}]=V\Lambda V^{\top}$, then $\operatorname{\mathbb{C}ov}[\hat{\beta}_{\rm MF}^{A},\hat{\beta}_{\rm MF}^{A}]$ has eigendecomposition $\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}]=\tilde{V}\Lambda\tilde{V}^{\top}$ where $\tilde{V}=(C_{XX})^{-1}V$. From Corollary 3.7, each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{C}_{XY}^{A},\hat{C}_{XY}^{A}]$ is minimized, so each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{\beta}_{\rm MF}^{A},\hat{\beta}_{\rm MF}^{A}]$ is also minimized, and thus the generalized variance $\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{\beta}_{\rm MF}^{A},\hat{\beta}_{\rm MF}^{A}])$ is minimized. ∎

### v2

@@ 3.3 Optimality analysis for control variate coefficients
This section proves the optimality of the control variate coefficient definitions in (15) and (16). To simplify the exposition, the analytical results are proved for the bi-fidelity case where $K=2$ and we drop the $\cdot_{k}$ subscripts for $A$ and $\alpha$, but the generalization of the results to $K>2$ case is straightforward.
The goal is to make our multifidelity learned model robust to random realizations of the training data: we thus begin by relating the variance of learned model predictions, conditioned on the given input $z$, to the covariance of the estimator $\hat{c}_{XY}^{\rm MF}$.
@@
For $i=1,\ldots,d$, let $(\lambda_{i},v_{i})$ denote the eigenpairs of $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$. Then,
$\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}^{\rm MF})|Z=z]=\sum_{i=1}^{d}\lambda_{i}c_{i}\leq\|c\|_{2}\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]),$ (17)
where $c=[c_{1},\ldots,c_{d}]^{\top}\in\mathbb{R}^{d}$ is a constant vector dependent on $z$ but independent of the estimator $\hat{\beta}^{\rm MF}$.
@@
$\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}^{\rm MF})|Z=z]$ $=\tilde{x}^{\top}\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]\tilde{x}=\sum_{i=1}^{d}\lambda_{i}\left<\tilde{x},v_{i}\right>^{2}.$
The equality in (17) follows from the assignment $c_{i}=\left<\tilde{x},v_{i}\right>^{2}$. Then, note that $\sum_{i=1}^{d}\lambda_{i}c_{i}\leq\|\lambda\|_{2}\|c\|_{2}\leq\|\lambda\|_{1}\|c\|_{2}$. The inequality follows from the fact that $\|\lambda\|_{1}=\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}])$. ∎
To minimize the variance of the learned model predictions, we therefore seek to minimize the eigenvalues of $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$. Let $g^{(k)}(z)=x(z)f^{(k)}(z)$ as before and note that $\hat{c}_{XY}^{\rm MF}$ can be rewritten as a multifidelity estimator for the mean of $g^{(1)}(z)$:
$\hat{c}_{XY}^{\rm MF}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(1)}(z_{i})+A\left(\frac{1}{m_{2}}\sum_{i=1}^{m_{2}}g^{(2)}(z_{i})-\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}g^{(2)}(z_{i})\right),$
where $A\in\mathbb{R}^{d\times d}$. Let $C_{ij}\in\mathbb{R}^{d\times d}=\operatorname{\mathbb{C}ov}[g^{(i)}(Z),g^{(j)}(Z)]$. Then, the autocovariance of $\hat{c}_{XY}^{\rm MF}$ is given by
$\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]=\frac{1}{m_{1}}\Gamma_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(A\Gamma_{22}A^{\top}-\Gamma_{12}A^{\top}-A\Gamma_{21}\right),$ (18)
We now show that the optimal matrix coefficient (15) minimizes the eigenvalues of (18).
@@
The choice $A^{*}=\Gamma_{12}\Gamma_{22}^{-1}$ minimizes all eigenvalues of (18).
@@
Let $D=A-\Gamma_{12}\Gamma_{22}^{-1}$. Then, we can rewrite (18) as
$\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$ $=\frac{1}{m_{1}}\Gamma_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(D(\Gamma_{22})^{-1}D^{\top}-\Gamma_{12}(\Gamma_{22})^{-1}\Gamma_{21}\right)$
$=\left(\frac{1}{m_{1}}\Gamma_{11}-\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\Gamma_{12}(\Gamma_{22})^{-1}\Gamma_{21}\right)+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)D(\Gamma_{22})^{-1}D^{\top}.$
Denote the first parenthetical term by $S_{1}\equiv\left(\frac{1}{m_{1}}\Gamma_{11}-\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\Gamma_{12}(\Gamma_{22})^{-1}\Gamma_{21}\right)$, and note that this term is symmetric and positive definite: grouping $1/{m_{1}}$ terms yields the conditional covariance of $g^{(1)}$ given $g^{(2)}$ which must be positive semi-definite and the $1/{m_{2}}$ term is positive. The second term above is also symmetric positive definite because it has the form $BB^{\top}$. Invoking Weyl’s inequality on the eigenvalues of Hermitian matrices (see [34, Corollary 4.3.12]), we have
$\lambda_{i}(S_{1})\leq\lambda_{i}\left(S_{1}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)D(\Gamma_{22})^{-1}D^{\top}\right),\quad\text{for }i=1,2,\ldots,d,$
where $\lambda_{i}(H)$ denotes the $i$-th eigenvalue of the Hermitian matrix $H$. Thus, each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$ is minimized for the choice $D=0$, and the conclusion follows. ∎
Theorem 3.4.
Let $A^{*}$ be given by Lemma 3.3. Then,
  1. 1.
    the generalized variance of the estimator $\hat{c}_{XY}^{\rm MF}$ is minimized,
  2. 2.
    the generalized variance of the estimated regression coefficients $\hat{\beta}^{\rm MF}$ is also minimized, and
  3. 3.
Proof.
The first statement follows directly from Lemma 3.3 since minimizing all eigenvalues of a matrix also minimizes its trace. To show the second statement, note that
$\operatorname{\mathbb{C}ov}[\hat{\beta}^{\rm MF},\hat{\beta}^{\rm MF}]=(C_{XX})^{-1}\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}](C_{XX})^{-\top},$
so if $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$ has eigendecomposition $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]=V\Lambda V^{\top}$, then $\operatorname{\mathbb{C}ov}[\hat{\beta}^{\rm MF},\hat{\beta}^{\rm MF}]$ has eigendecomposition $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]=\tilde{V}\Lambda\tilde{V}^{\top}$ where $\tilde{V}=(C_{XX})^{-1}V$. From Lemma 3.3, each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$ is minimized, so each eigenvalue of $\operatorname{\mathbb{C}ov}[\hat{\beta}^{\rm MF},\hat{\beta}^{\rm MF}]$ is also minimized, and thus the generalized variance $\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{\beta}^{\rm MF},\hat{\beta}^{\rm MF}])$ is minimized. The third statement follows from the equality in Lemma 3.2 and Lemma 3.3. ∎
In the case where $A=\alpha I$, the covariance expression (18) becomes:
$\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}]$ $=\frac{1}{m_{1}}\Gamma_{11}+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(\alpha^{2}\Gamma_{22}-\alpha\Gamma_{12}-\alpha\Gamma_{21}\right),$ (19)
We now show that the optimal scalar coefficient in (16) minimizes the trace of (19).
Lemma 3.5.
The choice $\alpha^{*}=\mathsf{Tr}(\Gamma_{12})/\mathsf{Tr}(\Gamma_{22})$ minimizes the expression for $\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}])$ in (19).
Proof.
The trace of the covariance of (19) is given by (recall $\mathsf{Tr}(M)=\mathsf{Tr}(M^{\top})$):
$\mathsf{Tr}(\operatorname{\mathbb{C}ov}[\hat{c}_{XY}^{\rm MF},\hat{c}_{XY}^{\rm MF}])=\frac{1}{m_{1}}\mathsf{Tr}(\Gamma_{11})+\left(\frac{1}{m_{1}}-\frac{1}{m_{2}}\right)\left(\alpha^{2}\mathsf{Tr}(\Gamma_{22})-2\alpha\mathsf{Tr}(\Gamma_{12})\right).$
@@
$2\alpha\mathsf{Tr}(\Gamma_{22})=2\mathsf{Tr}(\Gamma_{21})\implies\alpha=\mathsf{Tr}(\Gamma_{21})/\mathsf{Tr}(\Gamma_{22}).$
@@
Let $\alpha^{*}$ be given by Lemma 3.5 and let $A=\alpha^{*}I$. Then, within all possible control variate coefficients of the form $A=\alpha I$,
@@
    the trace of the covariance of the estimator $\hat{c}_{XY}$ is minimized, and